OpenStudy (anonymous):

rewrite in vertex form y = x^2+14x+50

OpenStudy (anonymous):

The vertex formula is y=a(x-h)^2 a=1 and to find the vertex you have to use the formula -b/2a -14/2(1) = -7 and that is the x! Fill in -7 for all the x's in y= x^2+14x+50 and you get 1 so your vertex is (-7,1) now just fill in everything into vertex form! y=1(x+7)^2+1

OpenStudy (anonymous):

Oops I meant the vertex formula is y=a(x-h)^2+k :)

OpenStudy (anonymous):

you are a life saver!!!!

OpenStudy (anonymous):

so for the work after finding out the answer for x is -7 i would write the equation as -7^2+14(-7)+50? @Anecia23

OpenStudy (anonymous):

and y=1 (x+7)^2+1 is the finished answer in vertex form? @Anecia23

OpenStudy (anonymous):

Yup!! :D